OpenStudy (anonymous):

Solve: 4x + 16 + 40 = 160 54 26 34 46

OpenStudy (anonymous):

you first have to combine like terms to simplify the expression. 4x + 16 + 40 = 160 16 + 40 = 56

OpenStudy (anonymous):

So now your equation looks like 4x + 56 = 160

OpenStudy (anonymous):

4x +16 +40 = 160 4x = 160 -16 -40 4x = 104 x = 104/4 x = 26